OpenStudy (anonymous):

Help Please! Find y when x=15 if y=16 when x=30.

OpenStudy (pawanyadav):

We can't say anything..

OpenStudy (pawanyadav):

There should be some more information ...

OpenStudy (anonymous):

This is complete! This is Direct Variation! This is exactly what the question says!

OpenStudy (pawanyadav):

No, this is not

OpenStudy (anonymous):

It is..You want Information?????? SHEESH! LEARN DIRECT VARIATION and... If y varies directly with x, write an equation for the direct variation

OpenStudy (anonymous):

@Pawanyadav ur no help...

OpenStudy (pawanyadav):

Oh sorry Its like this y=kx 16=k*30 k=16/30 Again. When x=15 y=k(15) Plug k y=16/30*15=16/2=8

OpenStudy (anonymous):

but still the question is incomplete

OpenStudy (anonymous):

what if it is like this->y=kx+c

OpenStudy (pawanyadav):

Direct variation.. y=kx

OpenStudy (anonymous):

yeah i think its like that y=kx but it shuld be specified in the ques

OpenStudy (pawanyadav):

Yes, the complete question was asked in another post.